1) The Giant ‘Darwin’s bark spider’ (Caerostris darwini), from Madagascar spins huge capture webs that can...

1)

The Giant ‘Darwin’s bark spider’ (Caerostris darwini), from Madagascar spins huge capture webs that can even cross small streams (No one really knows how they do it).

A web strand 10 µ in Diameter that is 10 cm long will extend to 10.1 cm under a 0.85 gram load. What is Young’s Modulus in GPa?

2)

The typical break stress for a hair is about 250 MPa. Suppose we had a giant circular hair that was 1 inch (2.54 cm) in diameter and had this break stress (250 MPa)

Approximately how many pounds of force would it take to break it? (1 N = 0.225 lbs-force)

The typical break stress for a hair is about 250 MPa. Suppose we had a giant circular hair that was 1 inch (2.54 cm) in diameter and had this break stress (250 MPa)

Approximately how many pounds of force would it take to break it? (1 N = 0.225 lbs-force)
Answer #1

Da 10am a 10 x 106 2 a f = 0.85*8 mm 10 x 10 en 5x10 m, Area, AzTR² 3.14 & (5 x 10-672 = 78.5 x 10 m2 0.85 x 10 -3 9.8= 8.334- 2 27 P 2 1 2 250 MPa z 250 * 10 Pa D= 2.540m 2 2.54 107m. A - T&P= 3:4 (1024x10-32 MB 20644103 2 1.27810?m A 5.06 x 10-4 m?